# Phonon fluctuation diagnostics: Origin of charge order in AV3Sb5 kagome metals

### v1

@@ Abstract
Introduction – Kagome metals have been an emerging ground for exotic states of electronic matter [1, 2, 3]. Instabilities in kagome metals are hypothesized to originate in simplified theoretical models from an intricate interplay of topology, electron-electron interactions, and enhanced availability of states due to Van Hove singularity (VHS) at the Fermi level [4, 5, 6, 7, 8, 9, 10]. Yet, kagome metals are real materials featuring higher degree of complexity than what is encoded in few-bands tight-binding fermionic-only models, so that the identification of electronic degrees of freedom relevant for the various instabilities remains a major outstanding point in the field. Here, we identify the electronic degrees of freedom responsible for the lattice instabilities in kagome metals directly. We find that these instabilities, across a series of compounds, involve electronic states deep inside the Brillouin zone (BZ), i.e., away from high symmetry points. This ab-initio based finding challenges current understanding based on low-energy VHS scenarios, nesting, and sublattice interference for Fermi surface instabilities [4, 5, 6, 7, 8]. Capable to adequately tracking down electronic degrees of freedom far of the Fermi level, our findings shed new light on the understanding of recent angle-resolved photoemission spectroscopy (ARPES).
@@
The family of AV3Sb5 (A = K, Rb, Cs) kagome superconductors has been heavily investigated due to its interesting lattice, frustration, topology, quantum geometry, chiral ordering, time-reversal symmetry breaking, flat bands, and the occurrence of a  charge-density wave (CDW) [11, 12, 13]. In particular, the origin of the charge ordering instability and its connection to superconductivity are currently under intense discussion [14]. The members of the AV3Sb5 group show a CDW transition below 78 K–103 K, where the structure undergoes a $2\times 2\times 2(4)$ reconstruction [15, 16, 17, 14]. At the level of a single layer (2D), the compound is assumed to adopt a Star of David (SD) or an inverse Star of David (ISD) configuration. Experimentally, the 135 V-based kagome exhibit different 3D ground-state reconstructions, where SD and or ISD can be stacked with or without a $\pi$-shift [18, 19]. The rearrangement can be directly linked to the theoretically observed phonon instabilities around the M- and L-points. We will put a particular focus on $\bm{q}=\mathrm{L}$; this is because we find it to exhibit the largest imaginary phonon frequency (Fig. 1(c)), even though either combinations of M and L can in principle contribute to the $2\times 2\times 2$ reconstruction (see Supplementary Information (SI) for $\bm{q}=\mathrm{M}$ analysis).
@@
The hitherto common narrative with regard to the origin of the CDW mechanism triangulates between the central themes of electronic interactions, phonon modes and electron-phonon coupling, as well as properties of Fermiology such as the nature and Fermi level distance of VHS, Fermi surface nesting, and sublattice interference [4, 5, 6, 7, 8]. One scenario of choice derives the charge order instability from the large density of states of the VHS at the M- and L-points close to the Fermi level. The phonon instability vectors $\bm{q}=\mathrm{M}$ and $\bm{q}=\mathrm{L}$ link these high-symmetry points, suggesting a Peierls mechanism resulting in a $2\times 2$ instability [4]. This explanation suggests a correlation between charge order and VHS proximity to the Fermi level, which is challenged by kagome CsTi3Bi5. There, the energy of the VHS can be tuned to the Fermi level with doping, yet no CDW is observed [23]. Even more drastically, recent findings reveal an anticorrelation between the CDW order parameter and the distance of the VHS to the Fermi level. Setting aside subtleties in the ordering propensities of charge vs nematic order [24] and its substantiation in the distinction between ordinary and higher order VHS, it suggests that the electronic states at the VHS might not be the single driving mechanism for the CDW [25]. Another related scenario centers around more general (Fermi surface) nesting, where, instead of specific high-symmetry points, entire lines in the BZ are connected by the instability vectors [5]. As much as this rational idea could explain the origin of the CDW for an idealistic system with perfect nesting, it struggles to account for real band structures under pressure and strain, where the Fermi surface nesting becomes significantly worse yet the charge order instability persists largely unaffected. While beforementioned scenarios imply a leading role of electronic correlation effects with an at best largely coarse-grained modelling of electron-phonon coupling, the charge order instability could also be determined by a strong or $\bm{k}$-selective electron-phonon coupling [6].
@@
In the following, the phonon self-energy will be further investigated. To disentangle the electronic response and matrix elements, we consider the momentum-resolved phonon self-energy $\Pi_{\bm{q}\mu\nu\bm{k}mn}$ as a product of the electron-phonon coupling $g^{2}_{\bm{q}\mu\nu\bm{k}mn}$ and the electronic susceptibility $\chi^{0}_{\bm{q}\bm{k}mn}$. This allows to distinguish the different contributions of the electronic density response to atomic displacements and enables us to identify the electronic processes behind the phonon softening associated with the origin of the CDW. We representatively discuss KV3Sb5 and refer to the SI for the other compounds. The relevant formula are:
@@
We analyze the phonon mode at $\bm{q}=\mathrm{L}$ where the highest imaginary frequency can be found, which is indicative of the instability in the system (phonon branch indices $\nu=\mu=\text{soft mode}$). The corresponding k-resolved electron-phonon coupling is obtained with ab-initio calculations and depicted in Fig. 4(a) for $k_{z}=0$ (for computational details, see Methods). We find the strongest electron-phonon coupling in the area between the M-points connected by the in-plane component of the chosen q vector, whereas not at the M-points themselves. To illustrate the electron-phonon coupling coefficient relevant for the phonon self-energy, in Fig. 4(a) we sum over the band indices $m$ and $n$ close to the Fermi level, as only these contribute to the product with the susceptibility (see Eq. 3). The electronic susceptibility evaluated based on Eq. 5 from a Wannier model is shown in Fig. 4(b) together with the $k_{z}=0$ Fermi surface in solid red lines. Additionally, a Fermi surface shifted by the instability wave vector $\bm{q}=\mathrm{L}$ is depicted with dashed lines to visualize possible low-energy transitions. As expected, the strongest contribution to the electron susceptibility can be found around the M-point (see SI for band-resolved susceptibility).
@@
A similar behavior can be observed for various $k_{z}$ cuts and the RbV3Sb5 and CsV3Sb5 compounds (see SI). This suggests a common origin of the CDW for the V-135 kagome group, even though different $2\times 2\times 2$ reconstructions are observed among these compounds, depending on temperature and pressure.
@@
Acknowledgements – We gratefully acknowledge the Gauss Centre for Supercomputing e.V. (https://www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at Leibniz Supercomputing Centre (https://www.lrz.de). We are grateful for funding support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter ct.qmat (EXC 2147, Project ID 390858490), the University Allowance of the University of Bremen (EXC 2077, Project ID 390741603), FOR 5249 (QUAST, Project ID 449872909 (Project P5), EXC 2056 (Cluster of Excellence “CUI: Advanced Imaging of Matter”, Project No. 390715994) as well as through the Collaborative Research Center SFB 1170 ToCoTronics (Project ID 258499086).

### v2

@@ Abstract
Introduction – Kagome metals have been an emerging ground for exotic states of electronic matter [1, 2, 3]. Instabilities in kagome metals are hypothesized to originate in simplified theoretical models from an intricate interplay of topology, electron-electron interactions, and enhanced availability of states due to Van Hove singularity (VHS) at the Fermi level [4, 5, 6, 7, 8, 9, 10]. Yet, kagome metals are real materials featuring higher degree of complexity than what is encoded in few-bands tight-binding fermionic-only models, so that the identification of electronic degrees of freedom relevant for the various instabilities remains a major outstanding point in the field. Here, we identify the electronic degrees of freedom responsible for the lattice instabilities in kagome metals directly. We find that these instabilities, across a series of compounds, involve electronic states deep inside the Brillouin zone (BZ), i.e., away from high symmetry points. This ab-initio based finding challenges current understanding based on low-energy VHS scenarios, nesting, and sublattice interference for Fermi surface instabilities [4, 5, 6, 7, 8, 11, 12]. Capable to adequately tracking down electronic degrees of freedom far of the Fermi level, our findings shed new light on the understanding of recent angle-resolved photoemission spectroscopy (ARPES).
@@
The family of AV3Sb5 (A = K, Rb, Cs) kagome superconductors has been heavily investigated due to its interesting lattice, frustration, topology, quantum geometry, chiral ordering, time-reversal symmetry breaking, flat bands, and the occurrence of a  charge-density wave (CDW) [13, 14, 15, 16]. In particular, the origin of the charge ordering instability and its connection to superconductivity are currently under intense discussion [17]. The members of the AV3Sb5 group show a CDW transition below 78 K–103 K, where the structure undergoes a $2\times 2\times 2(4)$ reconstruction [18, 19, 20, 17]. At the level of a single layer (2D), the compound is assumed to adopt a Star of David (SD) or an inverse Star of David (ISD) configuration. Experimentally, the 135 V-based kagome exhibit different 3D ground-state reconstructions, where SD and or ISD can be stacked with or without a $\pi$-shift [21, 22]. The rearrangement can be directly linked to the theoretically observed phonon instabilities around the M- and L-points. We will put a particular focus on $\bm{q}=\mathrm{L}$; this is because we find it to exhibit the largest imaginary phonon frequency (Fig. 1(c)), even though either combinations of M and L can in principle contribute to the $2\times 2\times 2$ reconstruction (see Supplementary Information (SI) for $\bm{q}=\mathrm{M}$ analysis [23]).
@@
The hitherto common narrative with regard to the origin of the CDW mechanism triangulates between the central themes of electronic interactions, phonon modes and electron-phonon coupling, as well as properties of Fermiology such as the nature and Fermi level distance of VHS, Fermi surface nesting, and sublattice interference [4, 5, 6, 7, 8, 11, 12]. One scenario of choice derives the charge order instability from the large density of states of the VHS at the M- and L-points close to the Fermi level. The phonon instability vectors $\bm{q}=\mathrm{M}$ and $\bm{q}=\mathrm{L}$ link these high-symmetry points, suggesting a Peierls mechanism resulting in a $2\times 2$ instability [4]. This explanation suggests a correlation between charge order and VHS proximity to the Fermi level, which is challenged by kagome CsTi3Bi5. There, the energy of the VHS can be tuned to the Fermi level with doping, yet no CDW is observed [27]. Even more drastically, recent findings reveal an anticorrelation between the CDW order parameter and the distance of the VHS to the Fermi level. Setting aside subtleties in the ordering propensities of charge vs nematic order [28] and its substantiation in the distinction between ordinary and higher order VHS, it suggests that the electronic states at the VHS might not be the single driving mechanism for the CDW [29]. Another related scenario centers around more general (Fermi surface) nesting, where, instead of specific high-symmetry points, entire lines in the BZ are connected by the instability vectors [5]. As much as this rational idea could explain the origin of the CDW for an idealistic system with perfect nesting, it struggles to account for real band structures under pressure and strain, where the Fermi surface nesting becomes significantly worse yet the charge order instability persists largely unaffected. While beforementioned scenarios imply a leading role of electronic correlation effects with an at best largely coarse-grained modelling of electron-phonon coupling, the charge order instability could also be determined by a strong or $\bm{k}$-selective electron-phonon coupling [6].
@@
In the following, the phonon self-energy will be further investigated. To disentangle the electronic response and matrix elements, we consider the momentum-resolved phonon self-energy $\Pi_{\bm{q}\mu\nu\bm{k}mn}$ as a product of the electron-phonon coupling $g^{2}_{\bm{q}\mu\nu\bm{k}mn}$ and the electronic susceptibility $\chi^{0}_{\bm{q}\bm{k}mn}$. This allows to distinguish the different contributions of the electronic density response to atomic displacements and enables us to identify the electronic processes behind the phonon softening associated with the origin of the CDW. We representatively discuss KV3Sb5 and refer to the SI for the other compounds [23]. The relevant formula are:
@@
We analyze the phonon mode at $\bm{q}=\mathrm{L}$ where the highest imaginary frequency can be found, which is indicative of the instability in the system (phonon branch indices $\nu=\mu=\text{soft mode}$). The corresponding k-resolved electron-phonon coupling is obtained with ab-initio calculations and depicted in Fig. 4(a) for $k_{z}=0$ (for computational details, see Methods). We find the strongest electron-phonon coupling in the area between the M-points connected by the in-plane component of the chosen q vector, whereas not at the M-points themselves. To illustrate the electron-phonon coupling coefficient relevant for the phonon self-energy, in Fig. 4(a) we sum over the band indices $m$ and $n$ close to the Fermi level, as only these contribute to the product with the susceptibility (see Eq. 3). The electronic susceptibility evaluated based on Eq. 5 from a Wannier model is shown in Fig. 4(b) together with the $k_{z}=0$ Fermi surface in solid red lines. Additionally, a Fermi surface shifted by the instability wave vector $\bm{q}=\mathrm{L}$ is depicted with dashed lines to visualize possible low-energy transitions. As expected, the strongest contribution to the electron susceptibility can be found around the M-point (see SI for band-resolved susceptibility [23]).
@@
A similar behavior can be observed for various $k_{z}$ cuts and the RbV3Sb5 and CsV3Sb5 compounds (see SI [23]). This suggests a common origin of the CDW for the V-135 kagome group, even though different $2\times 2\times 2$ reconstructions are observed among these compounds, depending on temperature and pressure.
@@
Acknowledgements – We gratefully acknowledge the Gauss Centre for Supercomputing e.V. (https://www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at Leibniz Supercomputing Centre (https://www.lrz.de). We are grateful for funding support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter ct.qmat (EXC 2147, Project ID 390858490), the University Allowance of the University of Bremen (EXC 2077, Project ID 390741603), FOR 5249 (QUAST, Project ID 449872909 (Project P3 and P5), EXC 2056 (Cluster of Excellence “CUI: Advanced Imaging of Matter”, Project No. 390715994) as well as through the Collaborative Research Center SFB 1170 ToCoTronics (Project ID 258499086).
Author contributions – S.E., J.B. and A.S. performed ab-initio calculations and carried out the phonon fluctuation diagnostics analysis. D.O., M.K. and R.C. conducted the ARPES experiments. T.W. and G.S. conceived and supervised this project. All authors contributed to the data interpretation and jointly wrote the manuscript.
Data availability – The data that support the findings of this letter are openly available [47, 48].